OpenStudy (needhelpfsat1234):

one person cleans floors in 4 hours another cleans them in 7. how long would it take if they worked together

satellite73 (satellite73):

you want to do it real real fast?

satellite73 (satellite73):

you can write down the answer instantly, compute it second

satellite73 (satellite73):

or we can do it the long slow boring way your choice

OpenStudy (rayep12):

et x=amount of time it takes both persons to clean the floor working together Then both persons, working together, cleans the floor at the rate of 1/x of the floor per hour The first person cleans floors at the rate of 1/4 of the floor per hour The other person cleans floors at the rate of 1/7 of the floor per hour So, our equation to solve is: (1/4)+(1/7)=1/x multiply each term by 28x 7x+4x=28 11x=28 x=28/11=~~~~~2.55 hours CK----- I'll use the fractions to avoid round-off errors. In (28/11) hours, the first person cleans (28/11)*(1/4) =7/11 of the floor In (28/11) hours, the other person cleans (28/11)*(1/7)=4/11 of the floor (7/11)+(4/11)=1---- one cleaned floor that is Hope this help

satellite73 (satellite73):

quick way \[\frac{4\times 7}{4+7}\]

OpenStudy (rayep12):

yeah nice

OpenStudy (needhelpfsat1234):

quick way

OpenStudy (needhelpfsat1234):

sry left for a minute